# Parameter-Efficient Fine-Tuning for Continual Learning: A Neural Tangent Kernel Perspective

### v1

@@ 6.1 Benchmark Comparison
In this subsection, we evaluate the NTK-CL method against other leading approaches. To ensure a fair performance comparison, we fix random seeds from 0 to 4, ensuring consistent task segmentation for each run  [^5] In Appendix D, detailed procedures for modifying class order and the class orders for primary datasets are provided, enabling researchers to accurately replicate our task segmentation process and evaluate the impact of different class orders on model performance.. We utilize uniformly sourced pre-trained weights and maintain the optimal hyper-parameters from the open-source code without modifications. Performance metrics for major datasets using ImageNet-21K pre-trained weights and ImageNet-1K fine-tuned weights are presented in Tables II and III.
For primary datasets such as CIFAR100, ImageNet-R, and ImageNet-A, we assess our method against most contemporary approaches, excluding DAP [36] due to its flawed testing process, Hide-Prompt [77] which compresses and samples past data, and Dual-PGP [62] which requires specific instance counts. By controlling for confounding factors, our method consistently achieves state-of-the-art performance. The NTK-CL method exhibits a clear advantage in both incremental accuracy ($\bar{A}$) and final accuracy ($A_{T}$), with improvements ranging from 1% to 7% compared to methods such as EASE [100] and EvoPrompt [39]. This advantage is particularly significant on ImageNet-A, a dataset known for challenging traditional models. Our approach substantially enhances model generalization and demonstrates robustness in complex visual recognition tasks.
Additionally, performance on auxiliary datasets including DomainNet, Oxford Pets, EuroSAT, PlantVillage, VTAB, and Kvasir, as detailed in Tables IV and V, highlights the generalization and adaptability of NTK-CL across diverse domains. On these datasets, NTK-CL not only consistently delivers superior accuracy metrics but also exhibits reduced variance in performance, emphasizing its stability. Notably, on Oxford Pets, NTK-CL achieves incremental accuracy improvements ranging from 1.8% to 2.1% and final accuracy enhancements of up to 4.6% compared to EASE [100]. On the Kvasir dataset, NTK-CL outperforms competing methods, achieving the highest incremental accuracy improvements ranging from 6.7% to 9.0% and the highest final accuracy improvements ranging from 19.3% to 21.1%, showcasing its significant potential for medical applications. Across other datasets, NTK-CL consistently ranks as the best or the second-best, further affirming the method’s efficacy and versatility. In conclusion, NTK-CL establishes a new benchmark in PEFT-CL, demonstrating superior performance across a diverse array of datasets.
@@ 6.2 Ablation Study
To ensure rigorous alignment between theoretical frameworks and empirical validation, a comprehensive suite of ablation studies is executed using the CIFAR100 and ImageNet-R datasets. Experimental conditions are standardized by setting the random seed to 0, ensuring consistent task segmentations, and utilizing weights from ViT-B/16-IN21K to maintain model consistency. The ablation studies encompass various configurations, including the Subnetwork-1 Adaptation Module (S1), Subnetwork-2 Adaptation Module (S2), Hybrid Adaptation Module (Hybrid), Knowledge Retention (KR), Task-Feature Dissimilarity Loss (Dis), Orthogonality Loss (Orth), and Regularization Loss (Reg). Average task-related accuracies ($\bar{A}$) are assessed to critically analyze their individual contributions to the model’s overall performance, which is displayed in Table VI.
Initial evaluations focus on the merits and drawbacks of S1, S2, and Hybrid modules, along with the benefits of their joint optimization. Experimental results across various datasets consistently demonstrate that the Hybrid module outperforms both S1 and S2 modules by effectively leveraging their combined strengths. Joint optimization leads to significant performance enhancement, affirming the generalization gains linked to increased sample size, which inherently supports knowledge transfer and retention in PEFT-CL.
The impact of the KR module across various adaptation module configurations is systematically explored to elucidate its influence on model performance. Implementing the KR module with the S1, S2, and Hybrid adaptation modules demonstrates substantial improvements in task-related accuracies ($\bar{A}$). Specifically, integrating the KR module increases the $\bar{A}$ from 82.99% to 85.45% on CIFAR100 and from 69.62% to 70.45% on ImageNet-R for the S1 configuration. For the S2 configuration, the $\bar{A}$ enhancement is more pronounced, rising from 85.04% to 91.37% on CIFAR100 and from 68.93% to 77.77% on ImageNet-R. The Hybrid module also benefits significantly, with increases from 86.51% to 89.50% on CIFAR100 and from 71.93% to 77.50% on ImageNet-R. A combined configuration of all three adaptation modules further raises the $\bar{A}$ to 92.01% on CIFAR100 and 81.08% on ImageNet-R, underscoring the KR module’s effectiveness in fostering knowledge retention and substantially boosting model generalization across diverse settings. It is noteworthy that while the KR module significantly enhances the performance of subnetwork-2 adaptation features, their potential remains limited compared to the jointly optimized hybrid features.
Further, the impact of task-feature dissimilarity losses is assessed, demonstrating that integrating Dis into the configuration results in $\bar{A}$ of 93.32% on CIFAR100 and 82.55% on ImageNet-R, showing enhancements of 4.30% and 8.43%. Additionally, the inclusion of Orth alongside the three adaptation modules and the KR module leads to an $\bar{A}$ of 92.39% on CIFAR100 and 81.10% on ImageNet-R, reflecting gains of 3.26% and 6.53%, respectively. These findings further emphasize the significant role of task-feature dissimilarity losses in PEFT-CL.
Finally, incorporating Reg within the same configuration results in an $\bar{A}$ of 92.39% on CIFAR100 and an enhanced $\bar{A}$ of 81.42% on ImageNet-R, indicating improvements of 3.26% and 6.95%, respectively. When all components and strategies are integrated, the model achieves the highest $\bar{A}$ values of 93.72% on CIFAR100 and 82.85% on ImageNet-R, reflecting cumulative enhancements of 4.75% and 8.83%.
Therefore, as summarized above, through these detailed ablation studies, we have demonstrated our theoretical derivations from a practical standpoint and ensured that each aspect of our design is rational and effective.
ViT-21K
S1
S2
ViT-21K
S1
S2
@@
TABLE VII: Performance analysis in NTK-CL for different pre-trained weights of ViT-B/16. The bolded segments represent the optimal results, while the underlined segments represent suboptimal results.
@@
6.3 Hyper-parameters Adjustment
Next, we will experiment with variations of the hyper-parameters $\eta$, $\upsilon$, and $\lambda$ to explore the impact of $\mathcal{L}_{dis}$, $\mathcal{L}_{orth}$, and $\mathcal{L}_{reg}$ on PEFT-CL performance to varying degrees. All experiments use a fixed random seed of 0 to ensure fair comparison. The optimal hyper-parameter settings differ between datasets. For ImageNet-R, the optimal values are set at $\eta=0.2$, $\upsilon=0.0001$, and $\lambda=0.001$. Conversely, for CIFAR100, the optimal settings are $\eta=0.03$, $\upsilon=0.0001$, and $\lambda=0.001$. During these experiments, while one parameter is varied, the other hyper-parameters are held constant to isolate the effects of the variable under study. The detailed tuning experiments for these hyper-parameters on ImageNet-R and CIFAR100 are displayed in Fig. 5.
As illustrated in Fig. 5, variations in $\eta$, $\upsilon$, and $\lambda$ significantly impact the performance of the PEFT-CL model during continual training. Keeping the values for orthogonality and regularization tightly around 0.0001 and 0.001, respectively, is crucial. Deviations can lead to significant performance degradation for both current and previous tasks. This highlights the challenge that overly stringent orthogonality between task features and parameter regularization can undermine the model’s basic classification performance. How to balance these aspects and eliminate negative impacts is a key issue that needs to be addressed in future PEFT-CL research.
6.4 Visualisation
To visually demonstrate the information captured by the pre-trained ViT and processed through the Subnetwork-1 and Subnetwork-2 Adaptation Modules, we select a random image from Task 0. We extract three-dimensional embeddings using parameters from these modules at various training stages and employ the Deep Image Prior (DIP) technique [74] to reconstruct the image information based on these embeddings. The Hybrid Adaptation Module, which uses two-dimensional features from the CLS Token, does not support DIP reconstruction visualizations. However, the embeddings from the Subnetwork-1 and Subnetwork-2 Adaptation Modules, which serve as inputs to the Hybrid Adaptation Module, sufficiently illustrate the network’s learning and memory capabilities. This approach offers a clear visual representation of how each module processes and retains information across different tasks, highlighting the dynamic learning and generalization capabilities of the adaptation modules in our NTK-CL framework.
In Fig. 6 and Fig. 7, we provide a detailed display of the DIP visualization results for images from Task-0 on ImageNet-R and ImageNet-A. These visualizations reveal distinct differences in the information captured by the Subnetwork-1 and Subnetwork-2 Adaptation Modules. Specifically, the Subnetwork-2 Adaptation Module tends to focus more on the shapes and inherent features of the images, whereas the Subnetwork-1 Adaptation Module emphasizes color and fine details. This distinction underscores our design strategy of differentiating feature subspaces and provides optimal input for the Hybrid Adaptation Module. Additionally, the evolution from Task-0 to Task-9 within our NTK-CL framework demonstrates its capability to effectively retain knowledge from previous tasks, affirming the efficacy of our Knowledge Retention innovation in maintaining consistent performance across tasks. In any embedding space generated by our adaptation modules, the information is well-maintained and aligns closely with the original pre-trained knowledge.
To further elucidate the evolution and advantages of subnetwork-1 adaptation features, subnetwork-2 adaptation features, and hybrid adaptation features during continual training, we conduct detailed t-SNE experiments. Utilizing the original ViT-21K pre-trained weights as a baseline, we compare the features of samples from Task 0 at both Task-0 and Task-9 stages. As illustrated in Fig. 8, the subnetwork-1 adaptation features, subnetwork-2 adaptation features, and hybrid adaptation features all exhibit significantly enhanced discriminability compared to the features produced by the original ViT-21K pre-trained weights. Additionally, the discriminability of Task-0 samples is effectively maintained even at Task-9, which to a certain extent demonstrates the anti-forgetting capability of our framework. Notably, the hybrid adaptation features show superior discriminability relative to both subnetwork-1 adaptation features and subnetwork-2 adaptation features, affirming the effectiveness of the fusion component in our framework.

### v2

@@ 6.1 Benchmark Comparison
In this subsection, we evaluate the NTK-CL method against other leading methods. To ensure a fair performance comparison, we fix random seeds from 0 to 4, ensuring consistent task segmentation for each run  [^5] In Appendix F, detailed procedures for modifying class order and the class orders for primary datasets are provided, enabling researchers to accurately replicate our task segmentation process and evaluate the impact of different class orders on model performance.. We utilize uniformly sourced pre-trained weights and maintain the optimal hyper-parameters from the open-source code without modifications. Performance metrics for major datasets using ImageNet-21K pre-trained weights and ImageNet-1K fine-tuned weights are presented in Tables II and III.
For primary datasets such as CIFAR100, ImageNet-R, and ImageNet-A, we assess our method against most contemporary methods, excluding DAP [39] due to its flawed testing process, Hide-Prompt [82] which compresses and samples past data, and Dual-PGP [66] which requires specific instance counts. By controlling for confounding factors, our method consistently achieves state-of-the-art performance. The NTK-CL method exhibits a clear advantage in both incremental accuracy ($\bar{A}$) and final accuracy ($A_{T}$), with improvements ranging from 1% to 7% compared to methods such as EASE [109] and EvoPrompt [43]. This advantage is particularly significant on ImageNet-A, a dataset known for challenging traditional models. Our NTK-CL framework substantially enhances model generalization and demonstrates robustness in complex visual recognition tasks.
Additionally, performance on auxiliary datasets including DomainNet, Oxford Pets, EuroSAT, PlantVillage, VTAB, and Kvasir, as detailed in Tables IV and V, highlights the generalization and adaptability of NTK-CL across diverse domains. On these datasets, NTK-CL not only consistently delivers superior accuracy metrics but also exhibits reduced variance in performance, emphasizing its stability. Notably, on Oxford Pets, NTK-CL achieves incremental accuracy improvements ranging from 1.8% to 2.1% and final accuracy enhancements of up to 4.6% compared to EASE [109]. On the Kvasir dataset, NTK-CL outperforms competing methods, achieving the highest incremental accuracy improvements ranging from 6.7% to 9.0% and the highest final accuracy improvements ranging from 19.3% to 21.1%, showcasing its significant potential for medical applications. Across other datasets, NTK-CL consistently ranks as the best or the second-best, further affirming the method’s efficacy and versatility.
To underscore the versatility of our NTK-CL framework, we provide a detailed examination of its performance in few-shot and imbalanced settings within Appendix L. Our results unequivocally illustrate that the framework sustains high performance levels under these conditions, validating the effectiveness of generalization principles.
@@ 6.2 Ablation Study
To rigorously align theoretical constructs with empirical evidence, an extensive series of ablation studies are conducted utilizing the CIFAR100 and ImageNet-R datasets. All experiments adhere to standardized conditions: a fixed random seed of 0, consistent task segmentation, and the utilization of pre-trained ViT-B/16-IN21K weight to ensure model consistency. The ablation studies examine configurations involving the Subnetwork-1 Adaptation Module (S1), Subnetwork-2 Adaptation Module (S2), Hybrid Adaptation Module (Hybrid), Knowledge Retention (KR) mechanism, Task-Feature Dissimilarity Loss (Dis), Orthogonality Loss (Orth), and Regularization Loss (Reg). The average accuracy ($\bar{A}$) across tasks is evaluated to quantitatively assess the contributions of each component to the model’s overall performance, which is displayed in Table VI.
Preliminary analyses reveal that the Hybrid module surpasses the standalone S1 and S2 modules by synergistically combining their strengths. This synergy, when optimized jointly, significantly boosts performance, highlighting the benefits of increased sample diversity in promoting knowledge transfer and retention within the PEFT-CL scenario. The efficacy of the KR module, in conjunction with different adaptation configurations, is systematically investigated. Results indicate that integrating the KR module markedly improves $\bar{A}$ across all adaptation modules. Notably, the KR module elevates the $\bar{A}$ from 82.99% to 85.45% on CIFAR100 and from 69.62% to 70.45% on ImageNet-R for the S1 configuration. For S2 module, the improvement is more pronounced, increasing from 85.04% to 91.37% on CIFAR100 and from 68.93% to 77.77% on ImageNet-R. The Hybrid module also sees a significant boost, with $\bar{A}$ rising from 86.51% to 89.50% on CIFAR100 and from 71.93% to 77.50% on ImageNet-R. Combining all three adaptation modules with the KR module achieves the highest $\bar{A}$ of 92.01% on CIFAR100 and 81.08% on ImageNet-R, underscoring the KR module’s pivotal role in enhancing knowledge retention and model generalization. The introduction of task-feature dissimilarity loss (Dis) further enhances performance, achieving $\bar{A}$ of 93.32% on CIFAR100 and 82.55% on ImageNet-R, representing improvements of 4.30% and 8.43%, respectively. Incorporating orthogonality loss (Orth) alongside the adaptation modules and KR module yields an $\bar{A}$ of 92.39% on CIFAR100 and 81.10% on ImageNet-R, indicating gains of 3.26% and 6.53%. Adding regularization loss (Reg) to this configuration further refines performance, achieving an $\bar{A}$ of 92.39% on CIFAR100 and 81.42% on ImageNet-R, with improvements of 3.26% and 6.95%. Integrating all components and strategies culminates in peak $\bar{A}$ values of 93.72% on CIFAR100 and 82.85% on ImageNet-R, marking cumulative enhancements of 4.75% and 8.83%. In summary, these ablation studies validate our theoretical hypotheses, confirming the effectiveness of each design choice.
TABLE VII: Comparison of feature size expansion methods and their impact on the evolution of incremental top-1 accuracy. The bold segments denote optimal results, and the underlined segments indicate suboptimal outcomes.
Combinations Type Incremental Top-1 Accuracy
Task 1 Task 2 Task 3 Task 4 Task 5 Task 6 Task 7 Task 8 Task 9 Task 10
Original + Mixed-up [105] 98.50 93.40 93.12 92.80 91.90 89.40 88.23 87.95 87.24 87.09
Original + PuzzleMix [41] 98.50 93.75 93.07 92.92 91.58 89.39 88.48 88.16 87.19 87.08
Original + AutoAug [16] 98.20 94.15 93.28 93.13 91.72 90.19 88.82 88.48 87.50 87.24
Original + RandAug [17] 98.20 94.05 93.42 93.10 92.02 90.41 89.80 89.18 88.09 87.99
IA3 [53] + S1 96.00 65.04 63.00 61.35 60.77 60.65 60.32 59.56 58.35 53.06
IA3 [53] + S2 98.50 96.35 95.07 94.20 93.54 91.21 91.18 89.94 89.09 88.59
Compacter [58] + S1 97.90 95.85 94.73 94.30 93.48 91.63 91.23 90.89 90.06 89.48
Compacter [58] + S2 98.60 96.75 95.37 94.80 94.30 92.04 92.03 91.26 90.59 90.03
Side-Tuning [106] + S1 95.80 93.55 91.70 89.48 88.36 85.89 85.82 85.06 84.02 83.18
Side-Tuning [106] + S2 98.50 96.55 95.73 94.10 93.62 91.33 91.20 90.46 89.92 89.11
Side-Tuning [106] + Compacter [58] 98.00 96.10 95.17 94.78 93.96 91.74 91.28 90.85 90.08 89.25
S1 + S2 (Ours) 98.70 96.65 95.70 94.85 94.36 92.28 92.07 91.44 90.90 90.24
TABLE VIII: Comparison of feature fusion methods and their impact on the evolution of incremental top-1 accuracy. $Q$, $K$, and $V$ denote the query, key, and value in our feature fusion method. The structures of the various fusion methods are illustrated in Appendix J. The bold segments denote optimal results, and the underlined segments indicate suboptimal outcomes.
Fusion Type Incremental Top-1 Accuracy
Task 1 Task 2 Task 3 Task 4 Task 5 Task 6 Task 7 Task 8 Task 9 Task 10
MLP 98.90 96.35 94.70 93.98 93.88 91.22 90.90 89.75 88.83 88.49
VAE 98.60 96.25 95.23 94.62 93.92 91.49 91.42 90.52 89.81 89.00
RNN 98.80 96.45 95.07 94.88 94.00 91.63 91.29 90.45 89.97 89.17
Mamba 98.30 96.75 95.50 94.68 93.84 91.63 91.24 90.20 89.86 89.12
S1($K/V$) + S2 ($Q$) 98.70 96.55 95.57 94.98 94.10 91.73 91.54 90.91 90.48 90.05
S1($Q$) + S2 ($K/V$) 98.70 96.65 95.70 94.85 94.36 92.28 92.07 91.44 90.90 90.24
TABLE IX: Comparison of knowledge inheritance methods and their impact on the evolution of incremental top-1 accuracy. The bold segments denote optimal results, and the underlined segments indicate suboptimal outcomes.
Knowledge Inheritance Type Incremental Top-1 Accuracy
Task 1 Task 2 Task 3 Task 4 Task 5 Task 6 Task 7 Task 8 Task 9 Task 10
MoCo-v3 [13] 98.70 95.90 93.17 91.32 90.50 87.63 87.31 87.01 86.21 85.36
LAE [22] 98.70 96.05 93.37 91.48 90.44 87.68 87.17 86.92 86.12 85.40
EASE [109] 98.70 96.65 91.77 91.45 91.36 91.15 90.89 90.49 89.36 88.23
Adaptive EMA (Ours) 98.70 96.65 95.70 94.85 94.36 92.28 92.07 91.44 90.90 90.24
TABLE X: Comparison of orthogonality losses and their impact on the evolution of incremental top-1 accuracy. The bold segments denote optimal results, and the underlined segments indicate suboptimal outcomes.
Orthogonality Loss Type Incremental Top-1 Accuracy
Name Level Task 1 Task 2 Task 3 Task 4 Task 5 Task 6 Task 7 Task 8 Task 9 Task 10
$\mathcal{L}_{\text{Standard}}$ Class-level 98.70 96.60 95.80 94.62 94.42 92.08 92.01 91.28 90.43 89.98
$\mathcal{L}_{\text{Schmidt}}$ Class-level 98.70 96.50 95.63 94.62 94.04 91.68 91.46 90.60 89.67 89.10
$\mathcal{L}_{\text{orth}}$ (Ours) Task-level 98.70 96.65 95.70 94.85 94.36 92.28 92.07 91.44 90.90 90.24
TABLE XI: Comparison of regularization methods and their impact on the evolution of incremental top-1 accuracy. The bold segments denote optimal results, and the underlined segments indicate suboptimal outcomes.
Regularization Type Incremental Top-1 Accuracy
Task 1 Task 2 Task 3 Task 4 Task 5 Task 6 Task 7 Task 8 Task 9 Task 10
L1 97.60 96.00 94.97 94.02 93.68 91.47 91.17 90.41 89.70 88.76
Spectral 98.60 96.60 95.70 94.70 94.34 92.04 92.00 91.12 90.57 89.06
HiDe-Prompt [82] 98.70 96.55 95.53 94.70 94.22 92.02 91.89 91.15 90.72 90.11
L2 (Ours) 98.70 96.65 95.70 94.85 94.36 92.28 92.07 91.44 90.90 90.24
6.3 Hyper-parameters Adjustment
In our experimental setup, we systematically vary the hyper-parameters $\eta$, $\upsilon$, and $\lambda$ to investigate their influence on the PEFT-CL performance, specifically the contributions of $\mathcal{L}_{dis}$, $\mathcal{L}_{orth}$, and $\mathcal{L}_{reg}$. To ensure a fair comparison across different conditions, all experiments employ a fixed random seed of 0. Our ultimately adopted optimal hyper-parameters are dataset-specific, with ImageNet-R benefiting from $\eta=0.2$, $\upsilon=0.0001$, and $\lambda=0.001$, whereas CIFAR100 achieves best results with $\eta=0.03$, $\upsilon=0.0001$, and $\lambda=0.001$. In each experiment, we isolate the effect of a single hyper-parameter by holding the others constant. As depicted in Fig. 5, alterations in these hyper-parameters markedly affect the model’s performance during continual learning. Maintaining orthogonality and regularization parameters near 0.0001 and 0.001, respectively, is essential for optimal performance. Deviating from these values can precipitate substantial declines in performance for both new and previously learned tasks, underscoring the delicate balance required between enforcing orthogonality among task features and applying parameter regularization to preserve classification accuracy.
Moreover, to address more complex real-world scenarios and mitigate the challenges associated with manual hyper-parameter specification, we propose two automatic hyper-parameter search (AHPS) methods: Bayesian Optimization strategy and Dynamic Loss Scaling strategy, as elaborated in Appendix K. The former offers a solid theoretical foundation but requires additional computational resources for validation-based search, whereas the latter operates without incurring extra computational overhead and eliminates the need for manual tuning of balancing coefficients. Empirical evaluations presented in Fig. 5 demonstrate that both automated strategies achieve performance on par with labor-intensive manual tuning, thereby providing practical benefits and enhanced efficiency for real-world PEFT-CL applications.
6.4 Alternative Experiments
To rigorously evaluate and underscore the distinct advantages of the proposed components, a series of meticulously designed alternative experiments are conducted on the CIFAR100 dataset. Each experiment adheres to a stringent protocol, employing a fixed random seed of 0 to ensure reproducibility, while all experimental conditions and runtime environments are rigorously standardized to maintain consistency. A particular emphasis is placed on the incremental top-1 accuracy $A_{\tau}$ across tasks.
The alternative experiments of sample size expansion methods encompasses two primary paradigms. Firstly, image-level augmentation techniques are explored, with Mixed-up [105], PuzzleMix [41], AutoAug [16], and RandAug [17] serving as representative methods. In this context, only the S2 component is retained, given its demonstrated superiority over S1 module on the CIFAR100 dataset, leveraging it to facilitate PEFT and the feature fusion from paired images. Despite achieving an expansion of the sample size at the image level, the results presented in Table VII reveal that this paradigm’s continual performance remains inferior to the network-level feature size expansion achieved through PEFT combinations. Furthermore, this paradigm incurs a higher computational cost, necessitating dual passes through both the backbone and subnetworks, as opposed to the single pass through the backbone and dual passes through the subnetworks required by the PEFT combinations. The second paradigm involves the amalgamation of distinct PEFT techniques, specifically IA3 [53], Compacter [58], and Side-Tuning [106]. As evidenced in Table VII, other PEFT combinations do not surpass the efficacy of ours. This disparity can be attributed to the fact that S1 module and S2 module extract information from the same image, albeit in the channel and spatial dimensions, respectively. This targeted extraction yields feature subspaces that are more discriminative and less redundant compared to those generated by other PEFT combinations that capture dataset-level biases, thereby contributing to superior performance.
In addition, systematic alternative experiments have been carried out concerning feature fusion, knowledge inheritance, orthogonality loss, and regularization loss. Comprehensive findings are summarized in Table VIII, Table IX, Table X, and Table XI respectively. Table VIII elucidates that the MSA technique, rigorously substantiated by theorem 4, stands out as the preeminent strategy for feature fusion. This method not only attains superior generalization loss but also excels in continual performance when juxtaposed with alternative fusion methods. Regarding knowledge inheritance, Table IX highlights the distinct superiority of the proposed Adaptive EMA mechanism, which facilitates effective retention and integration of historical knowledge, markedly outperforming its counterparts. The analysis presented in Table X reveals that, within the PEFT-CL scenario, task-level orthogonality suffices, diverging from conventional CL settings where class-level orthogonality is deemed essential. Lastly, the insights derived from Table XI affirm the validity of the adopted L2 regularization adjustment in solving saddle point dynamics in PEFT-CL, as delineated in Eq. 32. These findings collectively underscore the specificity and efficacy of our proposed components across various research dimensions.
@@ 6.4 Alternative Experiments
6.5 Visualization
To provide a more intuitive human visual assessment of the information captured by the pre-trained ViT and processed through the S1 and S2 modules, we demonstrate that this information resides in entirely distinct representational spaces. In Appendix I, we employ the Deep Image Prior (DIP) technique [79] to reconstruct the image information at different task stages. In Fig. 8 and Fig. 9, we present a detailed visualization of the DIP results for images from Task-0 on ImageNet-R and ImageNet-A datasets, respectively. These visualizations reveal distinct differences in the information captured by the S1 and S2 modules. Specifically, the S2 module tends to focus more on the shapes and intrinsic features of the images, while the S1 module emphasizes color and fine details. This distinction underscores our design strategy of differentiating feature subspaces, thereby providing optimal input for the Hybrid Adaptation Module. Furthermore, the evolution from Task-0 to Task-9 within our NTK-CL framework demonstrates its capability to effectively retain knowledge from previous tasks. This confirms the efficacy of our Knowledge Retention innovation in maintaining consistent performance across tasks, even as new tasks are introduced. These visualizations not only clarify the model’s behavior but also validate the effectiveness of our framework in the PEFT-CL scenario.
To further elucidate the evolution and advantages of S1 adaptation features, S2 adaptation features, and hybrid adaptation features during continual training, we conduct detailed t-SNE experiments. Utilizing the original ViT-21K pre-trained weights as a baseline, we compare the features of samples from Task 0 at both Task-0 and Task-9 stages. As illustrated in Fig. 6, the S1 adaptation features, S2 adaptation features, and hybrid adaptation features all exhibit significantly enhanced discriminability compared to the features produced by the original ViT-21K pre-trained weights. Additionally, the discriminability of Task-0 samples is effectively maintained even at Task-9, which to a certain extent demonstrates the anti-forgetting capability of our framework. Notably, the hybrid adaptation features show superior discriminability relative to both S1 adaptation features and S2 adaptation features, affirming the effectiveness of the fusion component.
TABLE XII: Performance analysis in NTK-CL for different pre-trained weights of ViT-B/16. The bolded segments represent the optimal results, while the underlined segments represent suboptimal results.